# VisionDirector: Vision-Language Guided Closed-Loop Refinement for Generative Image Synthesis

### v2

@@ 1 Introduction
The field of visual content creation has advanced rapidly with the rise of diffusion models, which now enable a wide range of applications in image generation and editing. However, professional designers still struggle to achieve faithful results because real-world design briefs often consist of long paragraphs with multiple goals, combining global art direction with numerous local constraints on characters, lighting, typography, and logo placement. However, existing benchmarks such as DrawBench[saharia2022photorealistictexttoimagediffusionmodels], TIFA[hu2023tifa], and MagicBrush[zhang2023magicbrush] predominantly use simplified prompts with at most one or two explicit goals, thereby masking model brittleness in realistic, multi-objective scenarios.
@@
To expose this gap and better evaluate models’ performance in real-world production settings, we introduce LongGoalBench (LGBench), which consists of 2,000 tasks sourced from artist workflows and includes multiple real-world targets per prompt. Specifically, each text-to-image (T2I) prompt specifies an average of 18.0 goals across 200 categories and 418 subcategories, for example, nature landscapes and historical scene, while the image-to-image (I2I) split, based on real source photos, contains an average of 11.2 edit directives. We evaluate several state-of-the-art models, including Flux-Krea [flux1kreadev2025], Flux-Kontex [flux2024], and Qwen-Image [wu2025qwen], on our proposed LGBench. As shown in Table 2, even the leading generative models, such as Flux-Kontex and Qwen-Image, only achieves 55.9% and 71.8% finish score, respectively. And they fail even more frequently on local text and logo edits. These results reveal why current models often fall short in real-world production settings. Although they demonstrate strong aesthetic abilities and can effectively handle single-goal instructions, they struggle with long and multi-goal instructions. Hence, a natural question arises: how can we enhance their ability to process multi-goal instructions without compromising their aesthetic quality?
@@
To solve this question, we propose VisionDirector, a training-free, VLM-guided [yang2025visionthink, chu2024towards] controller that decomposes long prompts into structured goals, decides whether to solve them in one shot or via staged micro-edits, and verifies/rolls back every edit. Instead of reporting a single reward at the end of generation, VisionDirector inspects each iteration, maintains goal-level state, and emits short instructions to the underlying diffusion models. This design lets us keep high-quality pretrained editors intact while adding a reasoning layer on top. Nevertheless, the controller can still be refined: we further fine-tune the planner with Group Relative Policy Optimization (GRPO), letting the agent experience thousands of long-goal rollouts and quickly learn when to STOP, VERIFY, or EDIT. By applying our proposed VisionDirector framework in post-training, the model is able to generate the desired images more quickly, reducing the average number of generation rounds from 4.2 to 3.1, achieving about a 26% reduction.
@@ 2.2 Benchmark Construction
Qwen-Image (T2I)  [wu2025qwen] T2I 71.8 30.7 76.6 29.7 87.0 78.3 85.8 77.5
Flux-Krea (T2I)  [flux1kreadev2025] T2I 66.8 18.6 75.7 25.6 82.3 67.7 78.1 67.8
Flux-Dev (T2I)  [flux2024] T2I 40.0 1.0 32.2 0.8 55.4 58.2 62.9 55.1
Qwen-Edit+ (I2I)  [wu2025qwen] I2I 71.0 52.4 62.0 93.2 60.3 77.0 70.9 49.3
Qwen-Edit (I2I)  [wu2025qwen] I2I 65.3 40.1 56.7 66.4 60.1 74.6 69.3 54.1
Flux-Kontext (I2I)  [batifol2025flux] I2I 55.9 31.0 49.7 70.8 48.3 62.6 53.6 39.7
@@
Property LGBench DrawBench [saharia2022photorealistic] TIFA [hu2023tifa] GenEval [ghosh2023geneval] MagicBrush [zhang2023magicbrush] ImgEdit [ye2025imgedit]
@@ 2.4 Discussion and Motivation
LGBench is designed to evaluate multi-attribute alignment rather than simple prompt fidelity. In real-world creative production, tasks rarely involve the one-shot fulfillment of a single textual cue; instead, they require the coherent realization of multiple interdependent goals such as composition, lighting, color harmony, and textual consistency. However, existing evaluation suites offer limited insight into this regime. As summarized in Table 3, previous benchmarks such as DrawBench [saharia2022photorealistic], and TIFA [hu2023tifa] rely on short prompts and human preference studies. GenEval [ghosh2023geneval] and MagicBrush [zhang2023magicbrush] introduce a limited number of QA-style or short edit tasks but still emphasize subjective or low-granularity scoring. None provide large-scale, reproducible evaluations in which success depends on meeting 10–23 quantitative goals verified automatically.
@@
Show-o  [xie2024show] 0.95 0.52 0.49 0.82 0.11 0.28 0.53
Emu3-Gen [wang2024emu3] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
PixArt-$\alpha$  [chen2023pixart] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SD3 Medium [esser2024scaling] 0.98 0.74 0.63 0.67 0.34 0.36 0.62
FLUX.1 [Dev]  [flux2024] 0.98 0.81 0.74 0.79 0.22 0.45 0.66
SD3.5 Large  [esser2024scalingrectifiedflowtransformers] 0.98 0.89 0.73 0.83 0.34 0.47 0.71
JanusFlow  [ma2025janusflow] 0.97 0.59 0.45 0.83 0.53 0.42 0.63
Lumina-Image 2.0  [qin2025lumina] 1.00 0.87 0.67 0.93 0.34 0.62 0.73
Janus-Pro-7B  [chen2025janus] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
HiDream-I1-Full  [cai2025hidream] 1.00 0.98 0.79 0.91 0.60 0.72 0.83
GPT Image 1 [High]  [openai2024gpt4technicalreport] 0.99 0.92 0.85 0.92 0.75 0.61 0.84
Seedream 3.0  [gao2025seedream] 0.99 0.96 0.91 0.93 0.47 0.80 0.84
Qwen-Image  [wu2025qwen] 0.99 0.92 0.89 0.88 0.76 0.77 0.87
@@ 3.1 Overview
Real-world creative tasks often require iterative coordination rather than one-shot generation. VisionDirector instantiates a closed-loop “director” that plans, executes, and verifies visual edits under long, interdependent instructions. The system integrates three modular components: (1) a planner VLM (Qwen3-VL-8B [bai2025qwen3vltechnicalreport] ) that parses and schedules goals, (2) text-to-image and image-to-image editors (Qwen-Image, Qwen-Image-Edit) that perform visual updates, and (3) a verifier VLM that provides goal-level feedback. The planner owns the semantics: it decomposes the user instruction into structured goals, decides execution order, and supervises revisions until convergence. Editors are stateless executors that follow the planner’s directives, while the verifier evaluates results and signals when to stop. Because each module communicates through natural-language messages, components can be swapped or upgraded without retraining the planner, and the same control policy can later be fine-tuned through reinforcement learning (§3.3).
@@ 3.2 Training-Free Closed-Loop Control
Figure 4 illustrates the overall workflow. The training-free controller operates through eight deterministic stages inspired by Deepseek-R1-style [guo2025deepseek] reasoning but specialized for visual generation.
@@ 3.3 GRPO for Efficient Generation
While the training-free VisionDirector already enhances instruction following, it still requires multiple rounds of editing before convergence. Furthermore, VisionDirector is not only a training-free pipeline but can also be employed after training to further improve generation efficiency. Specifically, we post-train the planner VLM using Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], a reinforcement learning method that encourages concise, high-reward edit strategies.
@@ 4.1 Evaluation Setup
Benchmarks. We evaluate VisionDirector on two benchmarks that stress long-horizon instruction following. For text-to-image (T2I), we report accuracy on GenEval [ghosh2023geneval]. GenEval focuses on compositional prompts with multiple objects and attributes. For image-to-image (I2I) editing we adopt ImgEdit [ye2025imgedit], which covers nine edit primitives.
@@ 4.3 Results on Image-to-Image
MagicBrush  [zhang2023magicbrush] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.90
Instruct-Pix2Pix  [brooks2023instructpix2pix] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit  [jiang2025anyedit] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit  [zhao2024ultraedit] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
OmniGen  [xiao2025omnigen] 3.47 3.04 1.71 2.94 2.43 3.21 4.19 2.24 3.38 2.96
ICEdit  [zhang2025context] 3.58 3.39 1.73 3.15 2.93 3.08 3.84 2.04 3.68 3.05
Step1X-Edit  [liu2025step1x] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
BAGEL  [deng2505emerging] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
UniWorld-V1  [lin2025uniworld] 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
OmniGen2  [wu2025omnigen2] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 4.68 3.44
FLUX.1 Kontext [Pro]  [batifol2025flux] 4.25 4.15 2.35 4.56 3.57 4.26 4.57 3.68 4.63 4.00
GPT Image 1 [High]  [openai2024gpt4technicalreport] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.89 4.20
Qwen-Image-Edit  [wu2025qwen] 4.38 4.16 3.43 4.66 4.14 4.38 4.81 3.82 4.69 4.27
@@ 5.1 Why we need agent framework for generation?
Our proposed VisionDirector explicitly delegates semantic processing to a planner VLM [yang2024visionzip], which extracts goals, reasons about conflicts, and verifies each edit, while the diffusion backbones focus solely on rendering. Through this separation of concerns, the model preserves visual quality, generates detailed logs for debugging, and maintains a modular architecture. Moreover, a more capable editor or verifier can be substituted without retraining the planner.
@@ 6 Related Work
Diffusion generation and editing. Large diffusion models such as DDPM/Score models [ho2020denoising] and latent diffusion variants [rombach2021highresolution] have driven recent leaps in visual quality. Follow-up work tailors them for instruction-based editing, e.g., Instruct-Pix2Pix [brooks2023instructpix2pix] or the more recent unified pipelines (Step1X-Edit, BAGEL, UniWorld) summarized in Table 5. Beyond images, diffusion-based video generation [che2024gamegenxinteractiveopenworldgame] and editing [yu2025positioninteractivegenerativevideo, yu2025surveyinteractivegenerativevideo] further magnify these limitations, as temporal consistency and long-horizon instruction adherence are difficult to maintain with single-shot prompting. These systems excel at local photorealism but still rely on single-shot prompts; none reason explicitly about dozens of heterogeneous goals. VisionDirector keeps these high-quality editors intact and adds a director layer that decomposes and monitors long instructions.
Benchmarking long-form instructions. Multi-goal evaluation has been explored in compositional prompts (GenEval [ghosh2023geneval]) and dense textual QA (DPG [hu2024ella]), but these suites cover at most two goals per prompt and focus on T2I only. ImgEdit [ye2025imgedit] and GEdit [liu2025step1x] evaluate instruction-based editing, yet their directives remain short and lack goal-level metadata. LGBench complements these efforts by providing dual modalities, 29k annotated goals, and a verifier that outputs structured pass/fail tags, enabling agent-style analysis.

### v3

@@ 1 Introduction
The field of visual content creation has advanced rapidly with the rise of diffusion models, which now enable a wide range of applications in image generation and editing. However, professional designers still struggle to achieve faithful results because real-world design briefs often consist of long paragraphs with multiple goals, combining global art direction with numerous local constraints on characters, lighting, typography, and logo placement. However, existing benchmarks such as DrawBench[27], TIFA[17], and MagicBrush[40] predominantly use simplified prompts with at most one or two explicit goals, thereby masking model brittleness in realistic, multi-objective scenarios.
@@
To expose this gap and better evaluate models’ performance in real-world production settings, we introduce LongGoalBench (LGBench), which consists of 2,000 tasks sourced from artist workflows and includes multiple real-world targets per prompt. Specifically, each text-to-image (T2I) prompt specifies an average of 18.0 goals across 200 categories and 418 subcategories, for example, nature landscapes and historical scene, while the image-to-image (I2I) split, based on real source photos, contains an average of 11.2 edit directives. We evaluate several state-of-the-art models, including Flux-Krea [20], Flux-Kontex [19], and Qwen-Image [31], on our proposed LGBench. As shown in Table 2, even the leading generative models, such as Flux-Kontex and Qwen-Image, only achieves 55.9% and 71.8% finish score, respectively. And they fail even more frequently on local text and logo edits. These results reveal why current models often fall short in real-world production settings. Although they demonstrate strong aesthetic abilities and can effectively handle single-goal instructions, they struggle with long and multi-goal instructions. Hence, a natural question arises: how can we enhance their ability to process multi-goal instructions without compromising their aesthetic quality?
@@
To solve this question, we propose VisionDirector, a training-free, VLM-guided [36, 8] controller that decomposes long prompts into structured goals, decides whether to solve them in one shot or via staged micro-edits, and verifies/rolls back every edit. Instead of reporting a single reward at the end of generation, VisionDirector inspects each iteration, maintains goal-level state, and emits short instructions to the underlying diffusion models. This design lets us keep high-quality pretrained editors intact while adding a reasoning layer on top. Nevertheless, the controller can still be refined: we further fine-tune the planner with Group Relative Policy Optimization (GRPO), letting the agent experience thousands of long-goal rollouts and quickly learn when to STOP, VERIFY, or EDIT. By applying our proposed VisionDirector framework in post-training, the model is able to generate the desired images more quickly, reducing the average number of generation rounds from 4.2 to 3.1, achieving about a 26% reduction.
@@ 2.2 Benchmark Construction
Qwen-Image (T2I)  [31] T2I 71.8 30.7 76.6 29.7 87.0 78.3 85.8 77.5
Flux-Krea (T2I)  [20] T2I 66.8 18.6 75.7 25.6 82.3 67.7 78.1 67.8
Flux-Dev (T2I)  [19] T2I 40.0 1.0 32.2 0.8 55.4 58.2 62.9 55.1
Qwen-Edit+ (I2I)  [31] I2I 71.0 52.4 62.0 93.2 60.3 77.0 70.9 49.3
Qwen-Edit (I2I)  [31] I2I 65.3 40.1 56.7 66.4 60.1 74.6 69.3 54.1
Flux-Kontext (I2I)  [2] I2I 55.9 31.0 49.7 70.8 48.3 62.6 53.6 39.7
@@
Property LGBench DrawBench [28] TIFA [17] GenEval [13] MagicBrush [40] ImgEdit [37]
@@ 2.4 Discussion and Motivation
LGBench is designed to evaluate multi-attribute alignment rather than simple prompt fidelity. In real-world creative production, tasks rarely involve the one-shot fulfillment of a single textual cue; instead, they require the coherent realization of multiple interdependent goals such as composition, lighting, color harmony, and textual consistency. However, existing evaluation suites offer limited insight into this regime. As summarized in Table 3, previous benchmarks such as DrawBench [28], and TIFA [17] rely on short prompts and human preference studies. GenEval [13] and MagicBrush [40] introduce a limited number of QA-style or short edit tasks but still emphasize subjective or low-granularity scoring. None provide large-scale, reproducible evaluations in which success depends on meeting 10–23 quantitative goals verified automatically.
@@
Show-o  [34] 0.95 0.52 0.49 0.82 0.11 0.28 0.53
Emu3-Gen [30] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
PixArt-$\alpha$  [6] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SD3 Medium [10] 0.98 0.74 0.63 0.67 0.34 0.36 0.62
FLUX.1 [Dev]  [19] 0.98 0.81 0.74 0.79 0.22 0.45 0.66
SD3.5 Large  [11] 0.98 0.89 0.73 0.83 0.34 0.47 0.71
JanusFlow  [23] 0.97 0.59 0.45 0.83 0.53 0.42 0.63
Lumina-Image 2.0  [25] 1.00 0.87 0.67 0.93 0.34 0.62 0.73
Janus-Pro-7B  [7] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
HiDream-I1-Full  [4] 1.00 0.98 0.79 0.91 0.60 0.72 0.83
GPT Image 1 [High]  [24] 0.99 0.92 0.85 0.92 0.75 0.61 0.84
Seedream 3.0  [12] 0.99 0.96 0.91 0.93 0.47 0.80 0.84
Qwen-Image  [31] 0.99 0.92 0.89 0.88 0.76 0.77 0.87
@@ 3.1 Overview
Real-world creative tasks often require iterative coordination rather than one-shot generation. VisionDirector instantiates a closed-loop “director” that plans, executes, and verifies visual edits under long, interdependent instructions. The system integrates three modular components: (1) a planner VLM (Qwen3-VL-8B [1] ) that parses and schedules goals, (2) text-to-image and image-to-image editors (Qwen-Image, Qwen-Image-Edit) that perform visual updates, and (3) a verifier VLM that provides goal-level feedback. The planner owns the semantics: it decomposes the user instruction into structured goals, decides execution order, and supervises revisions until convergence. Editors are stateless executors that follow the planner’s directives, while the verifier evaluates results and signals when to stop. Because each module communicates through natural-language messages, components can be swapped or upgraded without retraining the planner, and the same control policy can later be fine-tuned through reinforcement learning (§3.3).
@@ 3.2 Training-Free Closed-Loop Control
Figure 4 illustrates the overall workflow. The training-free controller operates through eight deterministic stages inspired by Deepseek-R1-style [14] reasoning but specialized for visual generation.
@@ 3.3 GRPO for Efficient Generation
While the training-free VisionDirector already enhances instruction following, it still requires multiple rounds of editing before convergence. Furthermore, VisionDirector is not only a training-free pipeline but can also be employed after training to further improve generation efficiency. Specifically, we post-train the planner VLM using Group Relative Policy Optimization (GRPO) [29], a reinforcement learning method that encourages concise, high-reward edit strategies.
@@ 4.1 Evaluation Setup
Benchmarks. We evaluate VisionDirector on two benchmarks that stress long-horizon instruction following. For text-to-image (T2I), we report accuracy on GenEval [13]. GenEval focuses on compositional prompts with multiple objects and attributes. For image-to-image (I2I) editing we adopt ImgEdit [37], which covers nine edit primitives.
@@ 4.3 Results on Image-to-Image
MagicBrush  [40] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.90
Instruct-Pix2Pix  [3] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit  [18] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit  [42] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
OmniGen  [33] 3.47 3.04 1.71 2.94 2.43 3.21 4.19 2.24 3.38 2.96
ICEdit  [41] 3.58 3.39 1.73 3.15 2.93 3.08 3.84 2.04 3.68 3.05
Step1X-Edit  [22] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
BAGEL  [9] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
UniWorld-V1  [21] 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
OmniGen2  [32] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 4.68 3.44
FLUX.1 Kontext [Pro]  [2] 4.25 4.15 2.35 4.56 3.57 4.26 4.57 3.68 4.63 4.00
GPT Image 1 [High]  [24] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.89 4.20
Qwen-Image-Edit  [31] 4.38 4.16 3.43 4.66 4.14 4.38 4.81 3.82 4.69 4.27
@@ 5.1 Why we need agent framework for generation?
Our proposed VisionDirector explicitly delegates semantic processing to a planner VLM [35], which extracts goals, reasons about conflicts, and verifies each edit, while the diffusion backbones focus solely on rendering. Through this separation of concerns, the model preserves visual quality, generates detailed logs for debugging, and maintains a modular architecture. Moreover, a more capable editor or verifier can be substituted without retraining the planner.
@@ 6 Related Work
Diffusion generation and editing. Large diffusion models such as DDPM/Score models [15] and latent diffusion variants [26] have driven recent leaps in visual quality. Follow-up work tailors them for instruction-based editing, e.g., Instruct-Pix2Pix [3] or the more recent unified pipelines (Step1X-Edit, BAGEL, UniWorld) summarized in Table 5. Beyond images, diffusion-based video generation [5] and editing [39, 38] further magnify these limitations, as temporal consistency and long-horizon instruction adherence are difficult to maintain with single-shot prompting. These systems excel at local photorealism but still rely on single-shot prompts; none reason explicitly about dozens of heterogeneous goals. VisionDirector keeps these high-quality editors intact and adds a director layer that decomposes and monitors long instructions.
Benchmarking long-form instructions. Multi-goal evaluation has been explored in compositional prompts (GenEval [13]) and dense textual QA (DPG [16]), but these suites cover at most two goals per prompt and focus on T2I only. ImgEdit [37] and GEdit [22] evaluate instruction-based editing, yet their directives remain short and lack goal-level metadata. LGBench complements these efforts by providing dual modalities, 29k annotated goals, and a verifier that outputs structured pass/fail tags, enabling agent-style analysis.